OpenStudy (anonymous):

2t^2+5t+?=0 How can i get the 3rd term?

OpenStudy (astrophysics):

You can complete the square

OpenStudy (phi):

the third term could be anything (unless we have more info) if we assume this equation is in standard form, the third term would be a number (with no "t" variable) for example, it could be 0, 1, 2, or pi , etc

OpenStudy (phi):

if we know the expression is a "perfect square" then we would do as astro suggests.

OpenStudy (astrophysics):

So if you have \[x^2+ax=0\] we can do the following \[x^2+\frac{ 2ax }{ 2 } + \left( \frac{ a }{ 2 } \right)^2-\left( \frac{ a }{ 2 } \right)^2=0\] (complete the square)

OpenStudy (anonymous):

your correct!